# Constraints on Compact Dark Matter Population from Micro-lensing Effect of Gravitational Wave for the third-generation gravitational Wave Detector

### v1

@@ Abstract
Since the pioneering detection of gravitational wave (GW) from a binary black hole merger by the LIGO-Virgo collaboration, GW has become a powerful tool for probing astrophysics and cosmology. If compact dark matter (DM) candidates, e.g. primordial black holes, contribute a substantial fraction of the DM component across a broad mass spectrum, their gravitational influence would imprint distinctive micro-lensing signatures on long-wavelength GW signals. In this paper, we propose a systematic method ranging from simulating micro-lensing GWs to constraining population information of compact DM, combining future third-generation ground-based GW detector, i.e. Einstein Telescope, with a hierarchical Bayesian inference framework. This approach enables us to analyze micro-lensing signatures in GW events and probe the potential contribution of compact DM objects to the DM budget of universe. For a population with a power-law mass function, we demonstrate that detecting 8 micro-lensing GW events out of $10^{4}$ observed binary black holes would constrain the compact DM abundance to $\log_{10}(f_{\rm CO})=-2.76^{+0.16}_{-0.17}$.
@@ I Introduction
The cosmic abundance of dark matter (DM), estimated to comprise $\sim 26\%$ of the total energy density in the universe, aligns with precision cosmological measurements (e.g. Cosmic Microwave Background anisotropies). Yet, despite this macroscopic confirmation, fundamental questions persist regarding its micro-physical properties and distribution at small scales. Particle physics and cosmology theories predict a variety of compact DM candidates, including massive compact halo objects (MACHOs), primordial black holes (PBHs), axion mini-clusters, compact mini-halos, boson stars, and fermion stars. In particular, PBH is taken as a potential candidate of DM and it has been a field of great astrophysical interest (see Sasaki et al. [1], Green and Kavanagh [2], Carr et al. [3] for recent reviews). To probe such compact objects, gravitational lensing effect is a powerful probe to constrain the abundance of compact DM over a broad mass range [4].
@@
On September 14, 2015, the gravitational wave (GW) detectors Advanced LIGO [5, 6] and Virgo [7] observed the first GW signal from binary black hole (BBH), GW150914 [8]. In 2019, KAGRA’s baseline installation was finalized [9, 10, 11]. From then on, the LIGO, Virgo, and KAGRA network, which is referred to as the LVK Collaboration, detected approximately $\mathcal{O}(100)$ GW signals during the O1, O2, and O3 observing runs [12, 13, 14, 15]. As more and more GW events are detected, a revolutionary window into the universe was opened. GWs offer unprecedented opportunities to explore astrophysics and cosmology. Similar to electromagnetic waves, GWs can be lensed by massive object, such as galaxy, compact DM. However, unlike stellar light, GWs detected by ground-based observatories have much longer wavelengths, making wave optics crucial [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. This diffraction effect distorts the GW signal, creating waveform fringes. Since Jung and Shin [27] first proposed that the micro-lensing effect of GWs detected by advanced LIGO can probe the compact DM, the micro-lensing of GW has been extensively studied to constrain on the compact DM in follow-up researches [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. However, the most existing studies focus solely on constraining the upper limits of abundance of compact DM ($f_{\rm CO}\equiv\frac{\Omega_{\rm CO}}{\Omega_{\rm DM}}$) through null detection of micro-lensing GW observations [27, 28, 32, 30, 31, 33, 29, 34, 35]. In addition, the significant increase in the sensitivity of next-generation ground-based GW detectors, i.e. Einstein Telescope (ET) [40, 41] and Cosmic Explorer (CE) [42], will enable the detection of micro-lensing GW events, which would provide more valuable information into the properties of compact DM. In this work, we present a comprehensive framework, spanning from simulating micro-lensing GW signals to extracting population information of compact DM embedded in these signals. Furthermore, we provide quantitative projections for the detection capabilities of next-generation ground-based GW detectors i.e. ET.
@@
II Methodology
In this Section, we present the theoretical framework for micro-lensing effects in GW, and develop a hierarchical Bayesian inference (HBI) approach to extract population properties of compact DM (e.g., mass function, and abundance) from an ensemble of micro-lensing GW events.
II.1 Theoretical framework for micro-lensing of gravitational wave
A GW is described by a small perturbation $h_{\mu\nu}$ over the background metric $g_{\mu\nu}^{(0)}$. The amplitude $h$ of this perturbation $h_{\mu\nu}$ satisfies the wave equation as
@@ II.1 Theoretical framework for micro-lensing of gravitational wave
Given micro-lensing events near the deflector, the lensed GW equation admits a Fourier-space approximation of the form
@@
where $U$ is the lensing potential in the background metric $g_{\mu\nu}^{(0)}$, and $\bar{h}$ is Fourier-transformed amplitude of $h$. We can define the dimensionless amplification factor as
@@
where $\bar{h}_{\rm L}(f)$ and $\bar{h}_{0}(f)$ correspond to lensed and unlensed ($U=0$) signal, respectively. The solution of Eq. (3) in the thin lens approximation is given by
$F(f)=\frac{D_{\rm s}D_{\rm l}\theta_{\rm E}^{2}(1+z_{\rm l})f}{D_{\rm ls}i}\int d^{2}\vec{x}\leavevmode\nobreak\ \exp[2\pi ift_{\rm d}(\vec{x},\vec{y})],$ (4)
where $D_{\rm s}$, $D_{\rm l}$ and $D_{\rm ls}$ represent the angular diameter distance to the source, to the lens, and between the source and the lens, respectively. In addition, $\theta_{\rm E}$ is Einstein radius, and $\vec{y}$ is the impact position of source. $t_{\rm d}$ is the arrival time delay of the wave to the observer
$t_{d}(\vec{x},\vec{y})=\frac{D_{\rm s}D_{\rm l}\theta_{\rm E}^{2}(1+z_{\rm l})}{D_{\rm ls}}\bigg{[}\frac{1}{2}|\vec{x}-\vec{y}|^{2}-\psi(\vec{x})-\phi_{\rm m}(\vec{y})\bigg{]},$ (5)
@@
where $\omega=8\pi m_{\rm l}(1+z_{\rm l})f$ is a dimensionless parameter and $_{1}F_{1}$ is the confluent hypergeometric function. Obviously, the amplification factor of the point mass lens depends only on the redshifted mass of lens $M_{\rm l}^{z}\equiv(1+z_{\rm l})m_{\rm l}$ and the source position $y$.
II.2 Hierarchical Bayesian Inference
@@
For population hyperparameters of compact DM $\Phi=[\bm{p}_{\rm mf},f_{\rm CO}]$ and $N_{\rm obs}$ detections of micro-lensing BBHs events $d=[d_{1},...,d_{N_{\rm obs}}]$, the likelihood follows an in-homogeneous Poisson distribution without any measurement uncertainty and selection effect is
@@
where $\Omega_{\rm DM}$ and $\rho_{\rm c}$ are the current fractional contribution of DM to the total energy density and critical density of the universe. In Eq. (9), $\sigma(m_{\rm l},z_{\rm l},z_{\rm s})$ is the lensing cross section
@@
It should be emphasized that we take the impact parameter as $y_{0}=5$, because $y_{0}>5$ are difficult to be identified as micro-lensing signals [29, 31]. In addition, $p_{\rm pop}(m_{\rm l}|\Phi)$ is the normalized distribution of lens masses and written as
@@
$\begin{split}N_{\rm det}(\Phi)=&\int dm_{\rm l}\int dz_{\rm s}\int_{0}^{z_{\rm s}}dz_{\rm l}\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}\times\\ &\frac{d\chi(z_{\rm l})}{dz_{\rm l}}(1+z_{\rm l})^{2}\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})N_{\rm s}P_{\rm s}(z_{\rm s}),\end{split}$ (13)
where $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is the cross section that micro-lensing signal can be detected and depends on Eq. (30) via
$\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{\max}(\lambda)^{2}.$ (14)
@@
where $p(\Phi)$ is the prior distribution for population hyperparameters $\Phi$. In addition, $Z_{\mathcal{M}}$ is both the normalized factor and the Bayesian evidence for the population model $\mathcal{M}$. This normalized factor can be calculated from the integral of the numerator of Eq. (16) over $\Phi$, i.e.
$Z_{\mathcal{M}}=\int d\Phi\leavevmode\nobreak\ p(d|\Phi)p(\Phi).$ (17)
III Simulation
In this Section, we introduce the simulation of GWs and how to generate micro-lensing GW events.
III.1 Simulation for model of GW sources
Although understanding the origin of GWs of BBHs is still an open issue, we select BBH merger events as the GW sources for this investigation due to some reasons: These systems exhibit high signal-to-noise ratio (SNR), robust waveform templates, and high event rates. The redshift distribution of GW sources is drawn from the phenomenological fit $R_{\rm BBH}$ to the population synthesis rate
$R_{\rm BBH}(z_{\rm s})=R_{0}\int_{t_{\min}}^{t_{\max}}R_{\rm f}(t(z_{\rm s})-t_{\rm d})P(t_{\rm d})dt_{\rm d},$ (19)
where $t_{\min}=50\leavevmode\nobreak\ {\rm Myr}$, $t_{\max}$ is set to the Hubble time, $P(t_{\rm d})\propto t_{\rm d}^{-1}$ is the time delay distribution which is set the power-law, and $R_{\rm f}(t(z_{\rm s})-t_{\rm d})$ is the Star Formation Rate (SFR) as [48, 30]
$R_{\rm f}(z_{\rm s})=[1+(1+z_{\rm p})^{-\lambda-\kappa}]\frac{(1+z_{\rm s})^{\lambda}}{1+\bigg{(}\frac{1+z_{\rm s}}{1+z_{\rm p}}\bigg{)}^{\lambda+\kappa}},$ (20)
where we set $\lambda=2.6$, $\kappa=3.6$, and $z_{\rm p}=2.2$. We use a power-law mass distribution for the heavier black hole in the BBHs as [28, 29, 30, 31]
$P_{\rm BBH}(m_{\rm s})=Z_{\rm m}m_{\rm s}^{-\alpha}\mathcal{H}(m_{\rm s}-m_{\min})\mathcal{H}(m_{\max}-m_{\rm s}),$ (21)
The heavier black hole $m_{1}$ satisfies the mass distribution as Eq.(21) with $\alpha=2.35$ in the range of $m_{1}\in[5,100]\leavevmode\nobreak\ M_{\odot}$, and the mass of $m_{2}$ uniformly distributes in the interval $m_{1}/18,m_{1}]$  [^1] For other parameters in GW sources, such as the orbital inclination and the merging orbital phase, we adopt a uniform distribution. In the following discussion, we do not show them, but we all take their influence into consideration..
III.2 Simulation for micro-lensing events
To ensure the physical rationality and computational efficiency of micro-lensing GW events simulation, we have developed a five-step process:
  • Generating the detectable BBH events: Based on the redshift distribution and mass distribution in above subsection, we can simulate $10^{4}$ detectable unlensed BBH events with $\rm{SNR}\geq 8$ for ET. The SNR of the unlensed GW signal can be computed as
    where $S_{\rm n}(f)$ is the power spectral density for ET.
  • Generating the micro-lensing GW events: By assuming that the compact DM is distributed uniformly in comving volume, the probability that a GW source at $z_{\rm s}$ is lensed by an intervening compact DM object can be given by optical depth as
    $P_{\rm L}(z_{\rm s})=1-\exp(-\tau(z_{\rm s})).$ (24)
    where $y_{0}$ is the maximum impact parameter. $\tau(m_{\rm l},z_{\rm l},y)$ is the differential optical depth for the mass of lens $m_{\rm l}$, redshift of lens $z_{\rm l}$ and impact parameter $y$, given by
    where we set $f_{\rm CO}=10^{-2.5}$.
  • Generating the parameters of lens: When a GW event is identified as lensed one, the value of $z_{\rm l}$ is adopted from a probability distribution $P(z_{\rm l})$ given by Eq. (26)
    $P(z_{\rm l})=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{z_{\rm s}}dz_{\rm l}\leavevmode\nobreak\ \tau(m_{\rm l},z_{\rm l},y)}.$ (27)
    $P(y)=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{y_{0}}dy\leavevmode\nobreak\ \tau(m_{\rm l},z_{\rm l},y)}=\frac{2}{y_{0}^{2}}y.$ (28)
    We set $y_{0}=5$ as mentioned above. For extended mass distribution of lens mass $\psi(m_{\rm l},\bm{p}_{\rm mf})$ same as in Eq. (10), we assume the maximum value of the extended power-law mass function is $M_{\max}=10^{3}\leavevmode\nobreak\ M_{\odot}$ and the minimum value is $M_{\min}=10\leavevmode\nobreak\ M_{\odot}$. The mass of the lens satisfies the power-law distribution described as
  • Computing lensed criterion: After generating unlensed and lensed GW waveform with simulated parameters for sources and lens, we use the mismatch $\epsilon$ and $\rm{SNR_{L}}$ as criterion for selection of lensing GW following [25]
    $\epsilon\times{\rm{SNR_{L}}}^{2}>1,$ (30)
    where mismatch $\epsilon$ is expressed as
  • Extract parameters of micro-lensing GWs: We use the Bayes’ theorem to forecast the constraints of the parameters $m_{1},m_{2},d_{\rm L}(z_{\rm s}),M_{\rm l}^{z},y]$ from lensed GW signals. The likelihood for each GW signals can be written in the form of an inner product [39]
From $10^{4}$ simulated GW events of BBHs, we identify 8 detectable micro-lensing GW signals. Moreover, we present the posterior distribution of the parameters of these events in Figure 1. As shown in the right panel of Figure 1, we compare the true simulated values of the luminosity distance for GW source and the redshift mass of lens with their posterior-derived counterparts. These results demonstrate good agreement within $2\sigma$ confidence levels.
IV Results

### v2

@@ Abstract
Since the pioneering detection of gravitational wave (GW) from a binary black hole merger by the LIGO-Virgo collaboration, GW has become a powerful probe for astrophysics and cosmology. If compact dark matter (DM) candidates, e.g. primordial black holes, contribute a substantial fraction of the DM component across a broad mass range, they would yield distinctive micro-lensing signatures on GW signals. In this paper, based on the third-generation ground-based GW detector, i.e. Einstein Telescope, we propose to constrain population information of compact DM by simulating micro-lensing GWs and analyzing with the hierarchical Bayesian inference framework. For a population with a power-law mass function, we demonstrate that detections of several micro-lensing GW signals in $10^{4}$ binary black holes coalescence events would constrain the abundance of compact DM to $\sim 10^{-3}$. It suggests that searching for and identifying micro-lensing signatures in future detections could be complementary and helpful in constraining compact DM scenarios.
@@ I Introduction
The cosmic abundance of dark matter (DM) has been estimated by precision cosmological measurements, such as the analysis of Cosmic Microwave Background anisotropies, to be approximately $26\%$ of the total energy density of the universe [1]. Despite this macroscopic confirmation, fundamental questions persist regarding its microphysical properties and distribution at small scales. Various hypotheses for dark matter may imply the presence of a variety of compact DM candidates, including primordial black holes (PBHs) [2, 3, 4], mini-halos [5, 6], boson stars [7, 8], and fermion stars [9]. In particular, PBH is taken as a potential candidate of DM and it has been a field of great astrophysical interest (see Sasaki et al. [10], Green and Kavanagh [11], Carr et al. [12] for recent reviews). To probe such compact objects, gravitational lensing effect is a powerful probe to constrain the abundance of compact DM over a broad mass range [13].
@@
On September 14, 2015, the gravitational wave (GW) detectors Advanced LIGO [14, 15] and Virgo [16] observed the first GW signal from binary black hole (BBH), GW150914 [17]. In 2019, KAGRA’s baseline installation was finalized [18, 19, 20]. From then on, the LIGO, Virgo, and KAGRA network, which is referred to as the LVK Collaboration, detected approximately $\mathcal{O}(100)$ GW signals during the O1, O2, and O3 observing runs [21, 22, 23, 24]. As more and more GW events are detected, a revolutionary window into the universe was opened. GWs offer unprecedented opportunities to explore astrophysics and cosmology. Similar to electromagnetic waves, GWs can be lensed by massive objects, such as galaxies, compact DM. However, unlike stellar light, GWs detected by observatories have much longer wavelengths, making wave optics crucial [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. This diffraction effect distorts the GW signal, creating waveform fringes. Since Jung and Shin [39] first proposed that the micro-lensing effect of GWs detected by Advanced LIGO can probe the compact DM, the micro-lensing of GW has been extensively studied to constrain on the compact DM in follow-up researches [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51]. However, the most existing studies focus solely on constraining the upper limits of abundance of compact DM ($f_{\rm CO}\equiv\Omega_{\rm CO}/\Omega_{\rm DM}$) through null detection of micro-lensing GW observations [39, 40, 44, 42, 43, 45, 41, 46, 47]. In addition, the significant increase in the sensitivity of next-generation ground-based GW detectors, i.e. Einstein Telescope (ET) [52, 53] and Cosmic Explorer (CE) [54], will enable the detection of micro-lensing GW events, which would provide more valuable information about the properties of compact DM. In this work, we present a comprehensive framework, spanning from simulating micro-lensing GW signals to extracting population information of compact DM embedded in these signals. Furthermore, we provide quantitative projections for the detection capabilities of next-generation ground-based GW detectors i.e. ET.
@@
II Theoretical framework for micro-lensing of gravitational wave
To describe the lensing effect of GWs, we consider gravitational potential of the lens $U(\vec{r})$ in the FLRW background $g_{\mu\nu}^{(0)}$,
$ds^{2}=-(1+2U(\vec{r}))dt^{2}+a(t)^{2}(1-2U(\vec{r}))d\vec{r}^{2}.$ (1)
GW is described by a small perturbation $h_{\mu\nu}$ over the background metric $g_{\mu\nu}^{(0)}$. The amplitude $h$ of this perturbation $h_{\mu\nu}$ satisfies the wave equation as
@@ II Theoretical framework for micro-lensing of gravitational wave
As a result of lensing being approximable as a local event occurring near the lens object, we can neglect the time dependence of the scale factor and write the Helmholtz equation for $h$ in Fourier space as
@@
where $\bar{h}$ is Fourier-transformed amplitude of $h$. We can define the dimensionless amplification factor as
@@
where $\bar{h}_{\rm L}(f)$ and $\bar{h}_{0}(f)$ correspond to lensed and unlensed ($U=0$) signal, respectively. The solution of Eq. (4) in the thin-lens and point source approximation is given by [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]
$F(f)=\frac{D_{\rm s}\xi_{0}^{2}(1+z_{\rm l})f}{iD_{\rm ls}D_{\rm l}}\int d^{2}\vec{x}\penalty 10000\ \exp[2\pi ift_{\rm d}(\vec{x},\vec{y})],$ (5)
where $\xi_{0}\equiv\theta_{\rm E}D_{\rm l}$ is a normalization factor with Einstein radius $\theta_{\rm E}$ to match the characteristic scale of the lensing configuration, $D_{\rm l}$ and $D_{\rm ls}$ represent the angular diameter distance to the source, to the lens, and between the source and the lens, respectively. In addition, the dimensionless vectors $\vec{x}$ and $\vec{y}$ are defined as
$\begin{split}&\vec{x}\equiv\frac{\vec{\xi}}{\xi_{0}},\\ &\vec{y}\equiv\frac{D_{\rm l}}{D_{\rm s}\xi_{0}}\vec{\eta},\end{split}$ (6)
where $\vec{\xi}$ and $\vec{\eta}$ are the physical coordinates of the image in the lens plane and of the source in the source plane, respectively. $\Delta t_{\rm d}$ is the arrival time delay of the wave to the observer
$\begin{split}&\Delta t_{d}(\vec{x},\vec{y})=\frac{D_{\rm s}\xi_{0}^{2}(1+z_{\rm l})}{D_{\rm l}D_{\rm ls}}\times\\ &\bigg[\frac{1}{2}|\vec{x}-\vec{y}|^{2}-\psi(\vec{x})-\phi_{\rm m}(\vec{y})\bigg],\end{split}$ (7)
@@
where $\omega=8\pi m_{\rm l}(1+z_{\rm l})f$ is a dimensionless parameter and $_{1}F_{1}$ is the confluent hypergeometric function. Obviously, the amplification factor of the point mass lens only depends on the redshifted mass of lens $M_{\rm l}^{z}\equiv(1+z_{\rm l})m_{\rm l}$ and the source position $y$. The geometrical optics limit corresponds to the scenario where the GW wavelength is significantly smaller than the Schwarzschild radius of the lens. In this regime, the amplification factor is well approximated by a sum of discrete magnification contributions from multiple images as
$\begin{split}F(f)=|\mu_{+}|^{1/2}-i|\mu_{-}|^{1/2}\exp(i2\pi f\Delta t_{\rm d}),\end{split}$ (9)
where $\mu_{+}$ and $\mu_{-}$ are the magnifications of the two lensing images.
III Simulation
In this section, we introduce the simulation of GWs and how we generate micro-lensing GW events. To ensure the physical rationality and computational efficiency of micro-lensing GW events simulation, we have developed a five-step process:
  • Generating the detectable BBH events: Although understanding the astrophysical origin of GWs of BBHs is still an open issue, we select BBH merger events as the GW sources for this investigation due to these reasons: These systems exhibit high signal-to-noise ratio (SNR), robust waveform templates, and high event rates. In addition, most GW sources observed so far by LVK Collaboration came from BBH mergers. The redshift distribution of GW sources is drawn from the phenomenological fit $R_{\rm BBH}$ to the population synthesis rate
    $\begin{split}&R_{\rm BBH}(z_{\rm s})=R_{0}\times\\ &\int_{t_{\min}}^{t_{\max}}R_{\rm f}(t(z_{\rm s})-t_{\rm d})P(t_{\rm d})dt_{\rm d},\end{split}$ (11)
    where $t_{\min}=50\penalty 10000\ {\rm Myr}$, $t_{\max}$ is set to the Hubble time, $P(t_{\rm d})\propto t_{\rm d}^{-1}$ is the time delay distribution which is set the power-law, and $R_{\rm f}(t(z_{\rm s})-t_{\rm d})$ inherits the redshift dependence of Star Formation Rate (SFR) as [56, 42]
    $\begin{split}R_{\rm f}(z_{\rm s})=[1+(1+z_{\rm p})^{-\lambda-\kappa}]\times\\ \frac{(1+z_{\rm s})^{\lambda}}{1+\bigg(\frac{1+z_{\rm s}}{1+z_{\rm p}}\bigg)^{\lambda+\kappa}},\end{split}$ (12)
    if $z_{\rm p}$ is much larger than 0 and $\lambda+\kappa$ is positive, this model is expected to behave like $(1+z)^{\lambda}$ and $(1+z)^{-\kappa}$ at low and high redshifts respectively. In Eq. (12), we set $\lambda=2.6$, $\kappa=3.6$, and $z_{\rm p}=2.2$ [56]. For mass function of GW sources, we use a power-law mass distribution for the heavier black hole in the BBHs as [40, 41, 42, 43]
    $\begin{split}&P_{\rm BBH}(m_{\rm s})=Z_{\rm m}m_{\rm s}^{-\alpha}\times\\ &\mathcal{H}(m_{\rm s}-m_{\min})\mathcal{H}(m_{\max}-m_{\rm s}),\end{split}$ (13)
    The heavier black hole $m_{1}$ satisfies the mass distribution as Eq.(13) with $\alpha=2.35$ in the range of $m_{1}\in[5,100]\penalty 10000\ M_{\odot}$, and the mass of $m_{2}$ uniformly distributes in the interval $m_{1}/18,m_{1}]$. For other parameters in GW sources, such as the spin ($s_{1}=s_{2}=0)$ and eccentricity ($e=0$), we adopt a fixed value. In addition, we take orbital inclination ($\iota\in[0,\pi]$), the coalescence phase ($\phi_{\rm c}\in[0,2\pi]$), location angles ($\theta\in[0,2\pi]$, $\phi\in[-\pi/2,\pi/2]$), and the polarisation angle ($\psi\in[0,\pi]$) as uniform distribution in GW waveform. We employ Python module PyCBC to produce IMRPhenomD waveform of frequency-domain for unlensed GW signals $\bar{h}_{0}(f)$ [57].
    Based on the redshift distribution and mass distribution in above subsection, we can simulate $10^{4}$ detectable unlensed BBH events with $\rm{SNR}\geq 8$ for ET. The SNR of the unlensed GW signal can be computed as
    where $S_{\rm n}(f)$ is the power spectral density for ET.
  • Generating the micro-lensing GW events: By assuming that the compact DM is distributed uniformly in comoving volume, the probability that a GW source at $z_{\rm s}$ is lensed by an intervening compact DM object is given by optical depth
    $P_{\rm L}(z_{\rm s})=1-\exp(-\tau(z_{\rm s})).$ (16)
    where $y_{0}$ is the maximum impact parameter. $\tau(m_{\rm l},z_{\rm l},y)$ is the differential optical depth evaluated at the mass of lens $m_{\rm l}$, redshift of lens $z_{\rm l}$ and impact parameter $y$, given by
    where we set $f_{\rm CO}=10^{-2.5}$.
  • Generating the parameters of lens: When a GW event is identified as a lensed one, the value of $z_{\rm l}$ is adopted from a probability distribution $P(z_{\rm l})$ as given in Eq. (18)
    $P(z_{\rm l})=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{z_{\rm s}}dz_{\rm l}\penalty 10000\ \tau(m_{\rm l},z_{\rm l},y)}.$ (19)
    $P(y)=\frac{\tau(m_{\rm l},z_{\rm l},y)}{\int_{0}^{y_{0}}dy\penalty 10000\ \tau(m_{\rm l},z_{\rm l},y)}=\frac{2}{y_{0}^{2}}y.$ (20)
    We set $y_{0}=5$ as mentioned above. For extended lens mass distribution $\psi(m_{\rm l},\bm{p}_{\rm mf})$ which is the same as that in Eq. (28), we assume the maximum value of the extended power-law mass function is $M_{\max}=10^{3}\penalty 10000\ M_{\odot}$ and the minimum value is $M_{\min}=10\penalty 10000\ M_{\odot}$. The mass of the lens satisfies the power-law distribution described as
  • Computing lensed criterion: After generating unlensed and lensed GW waveform with simulated parameters for sources and lens, we use the mismatch $\epsilon$ and $\rm{SNR_{L}}$ as criteria for the selection of lensing GW following [36]
    $\epsilon\times{\rm{SNR_{L}}}^{2}>1,$ (22)
    where the mismatch $\epsilon$ is expressed as
  • Extract parameters of micro-lensing GWs: We use the Bayes’ theorem to forecast the constraints of the parameters $m_{1},m_{2},d_{\rm L}(z_{\rm s}),M_{\rm l}^{z},y]$ from lensed GW signals. The likelihood for each GW signals can be expressed as an inner product [51]
From $10^{4}$ simulated BBH GW events, we identify 8 detectable micro-lensing GW signals. Moreover, we present the posterior distribution of the parameters of these events in Figure 1. As shown in the right panel of Figure 1, we compare the true simulated values of the luminosity distance for GW source and the redshift mass of lens with their posterior-derived counterparts. These results demonstrate good agreement within $2\sigma$ confidence levels.
IV Hierarchical Bayesian Inference
@@
For population hyperparameters of compact DM $\Phi$ and $N_{\rm obs}$ detections of micro-lensing BBHs events $d=[d_{1},...,d_{N_{\rm obs}}]$ ($N_{\rm obs}=8$ from $10^{4}$ simulated GW events as the simulation result in Section III), the likelihood follows an in-homogeneous Poisson distribution without any measurement uncertainty and selection effect is
@@
where $\Omega_{\rm DM}$ and $\rho_{\rm c}$ are the current fractional contribution of DM to the total energy density and critical density of the universe. In Eq. (28), population hyperparameters of compact DM are defined as
$\Phi\equiv[\bm{p}_{\rm mf},f_{\rm CO}].$ (29)
and $\sigma(m_{\rm l},z_{\rm l},z_{\rm s})$ in Eq. (27) is the lensing cross section
@@
It should be emphasized that we take the impact parameter as $y_{0}=5$, because when $y_{0}>5$, it’s difficult to identify the micro-lensing signal [41, 43]. In addition, $p_{\rm pop}(m_{\rm l}|\Phi)$ is the normalized distribution of lens masses and written as
@@
$\begin{split}N_{\rm det}(\Phi)=&\int dm_{\rm l}\int dz_{\rm s}\int_{0}^{z_{\rm s}}dz_{\rm l}\frac{dn(m_{\rm l},\Phi)}{dm_{\rm l}}\times\\ &\frac{d\chi(z_{\rm l})}{dz_{\rm l}}(1+z_{\rm l})^{2}\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})\times\\ &N_{\rm s}P_{\rm s}(z_{\rm s}),\end{split}$ (32)
where $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is the cross section of the micro-lensing signal that can be detected depending on Eq. (22). The cross section $\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})$ is given by
$\sigma_{\rm det}(\lambda,m_{\rm l},z_{\rm l},z_{\rm s})=\frac{4\pi m_{\rm l}D_{\rm l}D_{\rm ls}}{D_{\rm s}}y_{\max}(\lambda)^{2},$ (33)
@@
where $p(\Phi)$ is the prior distribution for population hyperparameters $\Phi$. In addition, $Z_{\mathcal{M}}$ is both the normalized factor and the Bayesian evidence for the population model $\mathcal{M}$.
V Results